OpenStudy (bunnyzrockgirl):

Write an equation in point-slope form for the line with the given slope that contains the given point. 1. slope = 5; (1, 7) Write an equation in slope-intercept from for the line through the two points. 2. (0, 2); (2, 6)

OpenStudy (anonymous):

lets start with question n#1

OpenStudy (bunnyzrockgirl):

Ok! Thanks for helping!

OpenStudy (anonymous):

u have a slope and a point. use with formula y-y1=m(x-x1) m=5 x1=1 y1=7

OpenStudy (bunnyzrockgirl):

so its y - 7 = 5(x-1) ?

OpenStudy (anonymous):

yes now solve for y

OpenStudy (bunnyzrockgirl):

i dont need to solve it, i just need the equation! Thanks, ive been told the formula before but no one has explained it! Can you help with question #2 now?

OpenStudy (anonymous):

and question n#2 you have been given 2 points but no m(slope).. in order to find the slope use this